Warmed #3 boiler, blewdown and checked L.W.C.O. on #1 boiler
Test and log boiler and tower water results.
Met Joe Hamm, he found problem with yard tower webmaster and made repair.He could not find problem with webmaster for #3and #4 tower’s.
I put boiler chemical pumps back on relay control from water meter.
Adjusted pumps up.
Talked to ARCH envoriomental rep. George,he said the Veeder Root alarm ( sudden loss ) will be OK until he comes in on the 17th.
Dale Mclauglin is requesting someone re-engerize the refrigerator in Labor & Delivery AP #4. Shut off to allow the unit to defrost. See w.o. # 821147 (room is vacant)
M.E.R. 20 switched off of Plant supplied chilled water ~ now on DX cooling. With o.a.t. temperatures dropping, need for chilled water should diminish.
Chuck and I replaced fan motor for trayline freezer compressor
Reset RF VFD for AHU 1 at SHB.